OpenStudy (anonymous):

Use the graph below to answer this question: graph of line going through negative 1, 5 and 4, negative 5 Find the average rate of change for the given function from x = −1 to x = 4.

Mehek (mehek14):

do you have the graph?

OpenStudy (anonymous):

its this- graph of line going through negative 1, 5 and 4, negative 5

OpenStudy (anonymous):

thats all i can do i can't get the image of it sorry

Mehek (mehek14):

click attach image

Mehek (mehek14):

are the points on the line (-1,5) and (4,-5)?

OpenStudy (anonymous):

i don't have the image but the coordinates are

Mehek (mehek14):

(-1,5) and (4,-5)?

OpenStudy (anonymous):

yes those are correct what you said

Mehek (mehek14):

\(\dfrac{-5-5}{4-(-1)}=\dfrac{-10}{5}=-2\)

OpenStudy (anonymous):

ok i was stuck between -2 and 2 . thank you! but could i ask you one more question?

Mehek (mehek14):

yea go ahead

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Jack and Jill had 3 pails of water. Each trip up the hill, Jack and Jill obtained 3 more buckets of water. Which of the graphs below represents Jack and Jill's number of pails over time? graph of line going through 0, 2 and 1, 6 graph of line going through 0, 2 and 4, 3 graph of line going through 0, 4 and 2, 5 graph of line going through 0, 3 and 1, 6

OpenStudy (anonymous):

sorry can't provide the graph pictures

Mehek (mehek14):

should be the last one

Mehek (mehek14):

because they already had 3 so 0,3 then each trip they got 3 more so in the first tripe = 1,6

OpenStudy (anonymous):

yeah thats what i was thinking because it goes through 3 but i wasn't sure

OpenStudy (anonymous):

thanks again!

Mehek (mehek14):

if it says goes up by 3, look for the answer choices in which it's going up by 3

Mehek (mehek14):

yw :D

OpenStudy (anonymous):

ok thanks!